# Study of mass outflows from magnetized accretion disks around rotating black holes with thermal conduction

### v1

@@ III Results
Figure 1: Variation of mass outflow rate ($R_{\dot{\rm m}}$) as a function of conduction parameter ($\Upsilon_{\rm s}$) for different $\lambda_{\rm edge}$ values, considering inflow injected from $x_{\rm edge}=1000$. Panels (a) and (b) depict results for $a_{\rm k}=0.0$ and $0.99$, respectively. In (a), we set $\alpha_{\rm B}=0.005$, $\mathcal{E}{\rm edge}=4\times 10^{-4}$, and $\beta{\rm edge}=10^{4}$, with $\lambda_{\rm edge}$ = $3.37$ (circle), $3.40$ (square), and $3.43$ (asterisk). In (b), $\alpha_{\rm B}=0.001$, $\mathcal{E}{\rm edge}=8\times 10^{-4}$, and $\lambda{\rm edge}$ = $2.07$ (circle), $2.08$ (square), and $2.09$ (asterisk). See the text for details.
@@
We present our numerical results of mass outflow rate ($R_{\dot{m}}$) around both weakly rotating ($a_{\rm k}\rightarrow 0$) and rapidly rotating ($a_{\rm k}=0.99$) black holes. In Fig.1, we depict the variation of $R_{\dot{\rm m}}$ with $\Upsilon_{\rm s}$ for flows with a fixed outer boundary. In Fig.1a, we inject matter around non-rotating black hole ($a_{\rm k}=0$) from the outer edge of the disk at $x_{\rm edge}=1000$ with $\mathcal{E}_{\rm edge}=4.0\times 10^{-4}$, $\alpha_{\rm B}=0.005$ and $\beta_{\rm edge}=10^{4}$. The dotted, dashed, and solid lines connecting the circles (green), diamonds (blue), and asterisks (red) are obtained for $\lambda_{\rm edge}=3.37$, $3.40$, and $3.43$, respectively. We observe that for a given $\lambda_{\rm edge}$, $R_{\dot{\rm m}}$ increases with the increase of $\Upsilon_{\rm s}$, which is consistent with the prediction of Tanaka and Menou (2006); Rezgui et al. (2019). Indeed, for a set of inflow parameters, increasing $\Upsilon_{\rm s}$ boosts the thermal pressure that eventually pushes the shock front outwards, thereby inflating the size of PSC Singh and Das (2024). As a result, the jet launching area increases. In addition, increased thermal conduction raises the local energy ($\mathcal{E}$) of the inflowing matter (see Fig 6 in Appendix B), which in turn increases the flow energy across the shock front ($x_{\rm s}$). Overall, a higher $\Upsilon_{\rm s}$ enhances the jet energy, providing a stronger driving force that ejects matter at a higher velocity leading to an enhanced outflow rate ($R_{\dot{\rm m}}$). On contrary, for a fixed $\Upsilon_{\rm s}$, $R_{\dot{\rm m}}$ increases with $\lambda_{\rm edge}$. This finding is not surprising because higher inflow angular momentum yields stronger centrifugal barrier, causing the shock to form at a larger radius, and hence, the size of PSC increases. Consequently, more of the inflowing material is deflected at the enhanced jet base, leading to an increase in the outflow rate ($R_{\dot{\rm m}}$). Next, in Fig.1b, we choose $\mathcal{E}_{\rm edge}=8.0\times 10^{-4}$, $\alpha_{\rm B}=0.001$, $\beta_{\rm edge}=10^{4}$ and inject matter around rapidly rotating black hole ($a_{\rm k}=0.99$) with different $\lambda_{\rm edge}$. Circles (green), squares (blue) and asterisks (red) joined with dotted, dashed and solid lines represent results for $\lambda_{\rm edge}=2.07$, $2.08$, and $2.09$, respectively. In general, we observe a strong corelation between $R_{\dot{\rm m}}$ and $\Upsilon_{\rm s}$ when $\lambda_{\rm edge}$ held fixed. Moreover, similar to the case of a weakly rotating black hole, $R_{\dot{\rm m}}$ remains higher for larger $\lambda_{\rm edge}$ when $\Upsilon_{\rm s}$ is fixed. Interestingly, outflows are generated from the rapidly rotating black hole even when $\lambda_{\rm edge}$ is relatively smaller compared to that of a weakly rotating black hole. This occurs primarily due to the spin-orbit coupling present in the effective potential of the black hole spacetime, where the marginally stable angular momentum strongly anti-correlates with the black hole spin $a_{\rm k}$. Furthermore, we observe that for a fixed $\lambda_{\rm edge}$, mass loss no longer occurs beyond a limiting value of $\Upsilon_{\rm s}$ (say $\Upsilon^{\rm max}_{\rm s}$), as RHCs are not favorable in that regime. Moreover, $\Upsilon^{\rm max}_{\rm s}$ decreases with $\lambda_{\rm edge}$ irrespective to $a_{\rm k}$ values. It is important to note that $\Upsilon^{\rm max}_{\rm s}$ does not assume universal values, as it primarily depends on the inflow parameters.
@@
Figure 2: Variation of $R_{\dot{\rm m}}$ with $\Upsilon_{\rm s}$ for a set of inflow energy $\mathcal{E}_{\rm edge}$. In panel (a), we choose $a_{\rm k}=0.0$, and inject matter from $x_{\rm edge}=1000$ with $\lambda_{\rm edge}=3.37$ and $\beta_{\rm edge}=10^{4}$ and $\alpha_{\rm B}=0.005$. Circles, squares and asterisks joined with dotted (green), dashed (blue) and solid (red) lines denote results for $\mathcal{E}_{\rm edge}=0.0003,0.0004$, and $0.0005$, respectively. In panel (b), we set $a_{\rm k}=0.99$ and $\alpha_{\rm B}=0.001$, and fix $\lambda_{\rm edge}=2.07$, and $\beta_{\rm edge}=10^{4}$ at $x_{\rm edge}=1000$. Circles, squares, and asterisks connected with dotted (green), dashed (blue), and solid (red) lines represent result for $\mathcal{E}_{\rm edge}$ = $5\times 10^{-4}$, $6.5\times 10^{-4}$, and $8\times 10^{-4}$, respectively. See the text for details.
@@
Next, we put effort to understand the influence of inflow energy in determining the mass outflow rate in presence of thermal conduction. In doing so, we examine how $R_{\dot{\rm m}}$ varies with $\Upsilon_{\rm s}$ for varied inflow energies. Here, we inject inflowing matter from the outer boundary at $x_{\rm edge}=10^{3}$ with different energy values ($\mathcal{E}_{\rm edge}$). The obtained results are shown in Fig. 2. In Fig. 2a, we present the results for non-rotating black hole with inflow parameters set to $\alpha_{\rm B}=0.005$, $\lambda_{\rm edge}=3.37$, and $\beta_{\rm edge}=10^{4}$. Dotted, dashed, and solid lines joining circles (green), squares (blue) and asterisks (red) correspond to the results obtained for $\mathcal{E}_{\rm edge}=3\times 10^{-4}$, $4\times 10^{-4}$, and $5\times 10^{-4}$, respectively. Similarly, in Fig. 2b, we present the variation of $R_{\dot{\rm m}}$ with $\Upsilon_{\rm s}$ around rapidly rotating black hole of spin $a_{\rm k}=0.99$. Here, we choose $\alpha_{\rm B}=0.001$, $\lambda_{\rm edge}=2.07$, and $\beta_{\rm edge}=10^{4}$. Results obtained for $\mathcal{E}_{\rm edge}=5\times 10^{-4}$, $6.5\times 10^{-4}$, and $8\times 10^{-4}$ are plotted using circles, squares and asterisks joined with dotted, dashed and solid lines. In both panels, we find that for a fixed $\Upsilon_{\rm s}$, $R_{\dot{\rm m}}$ increases with $\mathcal{E}_{\rm edge}$. This is expected, as higher $\mathcal{E}_{\rm edge}$ leads to the higher jet driving force resulting in an increased outflow rate.
@@
Figure 3: Variation of $R_{\dot{\rm m}}$ as a function of $\Upsilon_{\rm s}$ for a set of plasma-$\beta_{\rm edge}$ parameter. Panel (a) corresponds to BH spin parameter $a_{\rm k}=0.0$. Here, we choose $\alpha_{\rm B}=0.005$ and the other input parameters fixed at $x_{\rm edge}=10^{3}$ are $\lambda_{\rm edge}=3.37$ and $\mathcal{E}_{\rm edge}=4\times 10^{-4}$. Circles, squares, and asterisks joined with dotted (green), dashed (blue), and solid (red) lines represent results obtained for $\beta_{\rm edge}=10^{3}$, $2\times 10^{3}$ and $10^{4}$, respectively. In panel (b), we set $a_{\rm k}=0.99$ and $\alpha_{\rm B}=0.001$, and inject matter with $\lambda_{\rm edgee}=2.07$, and $\mathcal{E}_{\rm edge}=8\times 10^{-4}$ from $x_{\rm edge}=10^{3}$. Circles, squares, and asterisks joined with dotted (green), dashed (blue), and solid (red) lines are for $\beta_{\rm edge}$ = $500$, $10^{3}$, and $10^{4}$, respectively. See the text for details.
@@
We continue our investigation on mass loss as it is intriguing to explore the role of magnetic fields in generating outflows from the accretion disk in the presence of thermal conduction. To this end, we estimate $R_{\dot{\rm m}}$ by varying the plasma $\beta$ of the inflowing matter. The obtained results are shown in Fig. 3, where the variation of $R_{\dot{\rm m}}$ is depicted as a function of $\Upsilon_{\rm s}$ for a set of $\beta_{\rm edge}$ values. In Fig. 3a, we present the results for a non-rotating black hole ($a_{\rm k}=0.0$), where the inflowing matter is injected from the outer edge of the disk at $x_{\rm edge}=10^{3}$ with $\alpha_{\rm B}=0.005$, $\lambda_{\rm edge}=3.37$, and $\mathcal{E}_{\rm edge}=4\times 10^{-4}$. The circles, squares, and asterisks connected by the dotted (green), dashed (blue), and solid (red) lines denote results obtained for $\beta_{\rm edge}=10^{3}$, $2\times 10^{3}$, and $10^{4}$, respectively. Similarly, in Fig. 3b, we set $a_{\rm k}=0.99$ and fix $\alpha_{\rm B}=0.001$, $\lambda_{\rm edge}=2.07$, and $\mathcal{E}_{\rm edge}=8\times 10^{-4}$ for inflowing matter injected from $x_{\rm edge}=10^{3}$. The circles, squares, and asterisks joined using dotted (green), dashed (blue), and solid (red) lines are for $\beta_{\rm edge}$ = $5\times 10^{2}$, $10^{3}$, and $10^{4}$, respectively. From the figure, it is evident that $R_{\dot{\rm m}}$ generally increases with $\Upsilon_{\rm s}$ regardless of the strength of the disk magnetic fields ($\beta_{\rm edge}$) for both non-rotating and rapidly rotating black hole cases. Moreover, for a fixed value of $\Upsilon_{\rm s}$, a higher $\beta_{\rm edge}$ leads to an increased mass loss from the magnetized disk. As the magnetic field strength increases, it enhances angular momentum transport, causing the shock to settle down at a smaller radius. Because of this, the inflow energy at PSC increases that strengthens the jet-driving leading to more mass loss. It is worth noting that for rapidly rotating black hole ($a_{\rm k}=0.99$), the influence of the plasma-$\beta$ parameter on $R_{\dot{\rm m}}$ is more pronounced at relatively lower $\Upsilon_{\rm s}$ values, while $R_{\dot{\rm m}}$ tends converge at limiting $\Upsilon_{\rm s}$ values.
@@
Finally, we put efforts to estimate the maximum mass outflow rate ($R^{\rm max}_{\dot{\rm m}}$) from the magnetized accretion disk. To achieve this, we compute $R_{\dot{\rm m}}^{\rm max}$ as a function of $\Upsilon_{\rm s}$ for two extreme cases of BH spin: a non-rotating black hole ($a_{\rm k}=0.0$) and a rapidly rotating black hole ($a_{\rm k}=0.99$). Here, we set $\alpha_{\rm B}=0.01$ and fix the plasma-$\beta$ at the inner critical point ($r_{\rm in}$) as $\beta_{\rm in}=100$, while allowing the energy ($\mathcal{E}_{\rm in}$) and angular momentum ($\lambda_{\rm in}$) at $x_{\rm in}$ to vary freely. The obtained results are presented in Fig. 4, where circles connected by solid (red) and dashed (blue) lines representing results for non-rotating ($a_{\rm k}=0.0$) and rapidly rotating ($a_{\rm k}=0.99$) black holes, respectively. Our findings clearly reveal that higher $a_{\rm k}$ drives stronger outflows regardless of the conduction parameter ($\Upsilon_{\rm s}$), which remain consistent with the results reported earlier in weak thermal conduction limit Jana and Das (2024). Interestingly, for a fixed $a_{\rm k}$, $R_{\dot{\rm m}}^{\rm max}$ initially increases with $\Upsilon_{\rm s}$, however eventually tends to saturate. Furthermore, we note that for a rapidly rotating BH with $a_{\rm k}=0.99$, $R_{\dot{\rm m}}^{\rm max}$ peaks at an impressive $24\%$. These results indeed suggest that accretion flows around rapidly spinning black holes with higher thermal conduction are significantly more susceptible to mass loss.

### v2

@@ III Results
Figure 1: Variation of mass outflow rate ($R_{\dot{\rm m}}$) as a function of conduction parameter ($\Upsilon_{\rm s}$) for different $\lambda_{\rm edge}$ values, considering inflow injected from $x_{\rm edge}=1000$. Panels (a) and (b) depict results for $a_{\rm k}=0.0$ and $0.99$, respectively. In (a), we set $\alpha_{\rm B}=0.001$, $\mathcal{E}{\rm edge}=2\times 10^{-4}$, and $\beta{\rm edge}=10^{5}$, with $\lambda_{\rm edge}$ = $3.20$ (circle), $3.23$ (diamond), and $3.26$ (asterisk). In (b), $\alpha_{\rm B}=0.001$, $\mathcal{E}_{\rm edge}=1.2\times 10^{-3}$, and $\lambda{\rm edge}$ = $2.02$ (circle), $2.04$ (diamond), and $2.06$ (asterisk). See the text for details.
@@
We present our numerical results of mass outflow rate ($R_{\dot{m}}$) around both weakly rotating ($a_{\rm k}\rightarrow 0$) and rapidly rotating ($a_{\rm k}=0.99$) black holes. In Fig.1, we depict the variation of $R_{\dot{\rm m}}$ with $\Upsilon_{\rm s}$ for flows with a fixed outer boundary. In Fig.1a, we inject matter around non-rotating black hole ($a_{\rm k}=0$) from the outer edge of the disk at $x_{\rm edge}=1000$ with $\mathcal{E}_{\rm edge}=2.0\times 10^{-4}$, $\alpha_{\rm B}=0.001$ and $\beta_{\rm edge}=10^{5}$. The dotted, dashed, and solid lines connecting the circles (green), diamonds (blue), and asterisks (red) are obtained for $\lambda_{\rm edge}=3.20$, $3.23$, and $3.26$, respectively. We observe that for a given $\lambda_{\rm edge}$, $R_{\dot{\rm m}}$ increases with the increase of $\Upsilon_{\rm s}$, which is consistent with the prediction of Tanaka and Menou (2006); Rezgui et al. (2019). Indeed, for a set of inflow parameters, increasing $\Upsilon_{\rm s}$ boosts the thermal pressure that eventually pushes the shock front outwards, thereby inflating the size of PSC Singh and Das (2024). As a result, the jet launching area increases. In addition, increased thermal conduction raises the local energy ($\mathcal{E}$) of the inflowing matter (see Fig 6 in Appendix B), which in turn increases the flow energy across the shock front ($x_{\rm s}$). Overall, a higher $\Upsilon_{\rm s}$ enhances the jet energy, providing a stronger driving force that ejects matter at a higher velocity leading to an enhanced outflow rate ($R_{\dot{\rm m}}$). On contrary, for a fixed $\Upsilon_{\rm s}$, $R_{\dot{\rm m}}$ increases with $\lambda_{\rm edge}$. This finding is not surprising because higher inflow angular momentum yields stronger centrifugal barrier, causing the shock to form at a larger radius, and hence, the size of PSC increases. Consequently, more of the inflowing material is deflected at the enhanced jet base, leading to an increase in the outflow rate ($R_{\dot{\rm m}}$). Next, in Fig.1b, we choose $\mathcal{E}_{\rm edge}=1.2\times 10^{-3}$, $\alpha_{\rm B}=0.001$, $\beta_{\rm edge}=10^{5}$ and inject matter around rapidly rotating black hole ($a_{\rm k}=0.99$) with different $\lambda_{\rm edge}$. Circles (green), diamonds (blue) and asterisks (red) joined with dotted, dashed and solid lines represent results for $\lambda_{\rm edge}=2.02$, $2.04$, and $2.06$, respectively. In general, we observe a strong correlation between $R_{\dot{\rm m}}$ and $\Upsilon_{\rm s}$ when $\lambda_{\rm edge}$ held fixed. Moreover, similar to the case of a weakly rotating black hole, $R_{\dot{\rm m}}$ remains higher for larger $\lambda_{\rm edge}$ when $\Upsilon_{\rm s}$ is fixed. Interestingly, outflows are generated from the rapidly rotating black hole even when $\lambda_{\rm edge}$ is relatively smaller compared to that of a weakly rotating black hole. This occurs primarily due to the spin-orbit coupling present in the effective potential of the black hole spacetime, where the marginally stable angular momentum strongly anti-correlates with the black hole spin $a_{\rm k}$. Furthermore, we observe that for a fixed $\lambda_{\rm edge}$, mass loss no longer occurs beyond a limiting value of $\Upsilon_{\rm s}$ (say $\Upsilon^{\rm max}_{\rm s}$), as RHCs are not favorable in that regime. Moreover, $\Upsilon^{\rm max}_{\rm s}$ decreases with $\lambda_{\rm edge}$ irrespective to $a_{\rm k}$ values. It is important to note that $\Upsilon^{\rm max}_{\rm s}$ does not assume universal values, as it primarily depends on the inflow parameters.
@@
Figure 2: Variation of $R_{\dot{\rm m}}$ with $\Upsilon_{\rm s}$ for a set of inflow energy $\mathcal{E}_{\rm edge}$. In panel (a), we choose $a_{\rm k}=0.0$, and inject matter from $x_{\rm edge}=1000$ with $\lambda_{\rm edge}=3.20$ and $\beta_{\rm edge}=10^{5}$ and $\alpha_{\rm B}=0.001$. Circles, diamonds and asterisks joined with dotted (green), dashed (blue) and solid (red) lines denote results for $\mathcal{E}_{\rm edge}=0.0002,0.0003$, and $0.0004$, respectively. In panel (b), we set $a_{\rm k}=0.99$ and $\alpha_{\rm B}=0.001$, and fix $\lambda_{\rm edge}=2.02$, and $\beta_{\rm edge}=10^{5}$ at $x_{\rm edge}=1000$. Circles, diamonds, and asterisks connected with dotted (green), dashed (blue), and solid (red) lines represent result for $\mathcal{E}_{\rm edge}$ = $8.0\times 10^{-4}$, $1.0\times 10^{-3}$, and $1.2\times 10^{-3}$, respectively. See the text for details.
@@
Next, we put effort to understand the influence of inflow energy in determining the mass outflow rate in presence of thermal conduction. In doing so, we examine how $R_{\dot{\rm m}}$ varies with $\Upsilon_{\rm s}$ for varied inflow energies. Here, we inject inflowing matter from the outer boundary at $x_{\rm edge}=10^{3}$ with different energy values ($\mathcal{E}_{\rm edge}$). The obtained results are shown in Fig. 2. In Fig. 2a, we present the results for non-rotating black hole with inflow parameters set to $\alpha_{\rm B}=0.001$, $\lambda_{\rm edge}=3.20$, and $\beta_{\rm edge}=10^{5}$. Dotted, dashed, and solid lines joining circles (green), diamonds (blue) and asterisks (red) correspond to the results obtained for $\mathcal{E}_{\rm edge}=2\times 10^{-4}$, $3\times 10^{-4}$, and $4\times 10^{-4}$, respectively. Similarly, in Fig. 2b, we present the variation of $R_{\dot{\rm m}}$ with $\Upsilon_{\rm s}$ around rapidly rotating black hole of spin $a_{\rm k}=0.99$. Here, we choose $\alpha_{\rm B}=0.001$, $\lambda_{\rm edge}=2.02$, and $\beta_{\rm edge}=10^{5}$. Results obtained for $\mathcal{E}_{\rm edge}=8.0\times 10^{-4}$, $1.0\times 10^{-3}$, and $1.2\times 10^{-3}$ are plotted using circles, diamonds and asterisks joined with dotted, dashed and solid lines. In both panels, we find that for a fixed $\Upsilon_{\rm s}$, $R_{\dot{\rm m}}$ increases with $\mathcal{E}_{\rm edge}$. This is expected, as higher $\mathcal{E}_{\rm edge}$ leads to the higher jet driving force resulting in an increased outflow rate.
@@
Figure 3: Variation of $R_{\dot{\rm m}}$ as a function of $\Upsilon_{\rm s}$ for a set of plasma-$\beta_{\rm edge}$ parameter. Panel (a) corresponds to BH spin parameter $a_{\rm k}=0.0$. Here, we choose $\alpha_{\rm B}=0.001$ and the other input parameters fixed at $x_{\rm edge}=1000$ are $\lambda_{\rm edge}=3.20$ and $\mathcal{E}_{\rm edge}=2\times 10^{-4}$. Circles, diamonds, and asterisks joined with dotted (green), dashed (blue), and solid (red) lines represent results obtained for $\beta_{\rm edge}=5\times 10^{3}$, $10^{4}$ and $10^{5}$, respectively. In panel (b), we set $a_{\rm k}=0.99$ and $\alpha_{\rm B}=0.001$, and inject matter with $\lambda_{\rm edgee}=2.02$, and $\mathcal{E}_{\rm edge}=1.2\times 10^{-3}$ from $x_{\rm edge}=10^{3}$. Circles, diamonds, and asterisks joined with dotted (green), dashed (blue), and solid (red) lines are for $\beta_{\rm edge}$ = $5\times 10^{3}$, $10^{4}$, and $10^{5}$, respectively. See the text for details.
@@
We continue our investigation on mass loss as it is intriguing to explore the role of magnetic fields in generating outflows from the accretion disk in the presence of thermal conduction. To this end, we estimate $R_{\dot{\rm m}}$ by varying the plasma $\beta$ of the inflowing matter. The obtained results are shown in Fig. 3, where the variation of $R_{\dot{\rm m}}$ is depicted as a function of $\Upsilon_{\rm s}$ for a set of $\beta_{\rm edge}$ values. In Fig. 3a, we present the results for a non-rotating black hole ($a_{\rm k}=0.0$), where the inflowing matter is injected from the outer edge of the disk at $x_{\rm edge}=1000$ with $\alpha_{\rm B}=0.001$, $\lambda_{\rm edge}=3.20$, and $\mathcal{E}_{\rm edge}=2\times 10^{-4}$. The circles, diamonds, and asterisks connected by the dotted (green), dashed (blue), and solid (red) lines denote results obtained for $\beta_{\rm edge}=5\times 10^{3}$, $10^{4}$, and $10^{5}$, respectively. Similarly, in Fig. 3b, we set $a_{\rm k}=0.99$ and fix $\alpha_{\rm B}=0.001$, $\lambda_{\rm edge}=2.02$, and $\mathcal{E}_{\rm edge}=1.2\times 10^{-3}$ for inflowing matter injected from $x_{\rm edge}=1000$. The circles, diamonds, and asterisks joined using dotted (green), dashed (blue), and solid (red) lines are for $\beta_{\rm edge}$ = $5\times 10^{3}$, $10^{4}$, and $10^{5}$, respectively. From the figure, it is evident that $R_{\dot{\rm m}}$ generally increases with $\Upsilon_{\rm s}$ regardless of the strength of the disk magnetic fields ($\beta_{\rm edge}$) for both non-rotating and rapidly rotating black hole cases. Moreover, for a fixed value of $\Upsilon_{\rm s}$, a higher $\beta_{\rm edge}$ leads to an increased mass loss from the magnetized disk. As the magnetic field strength increases, it enhances angular momentum transport, causing the shock to settle down at a smaller radius. Because of this, the inflow energy at PSC increases that strengthens the jet-driving leading to more mass loss. It is worth noting that for rapidly rotating black hole ($a_{\rm k}=0.99$), the influence of the plasma-$\beta$ parameter on $R_{\dot{\rm m}}$ is more pronounced at relatively lower $\Upsilon_{\rm s}$ values, while $R_{\dot{\rm m}}$ tends converge at limiting $\Upsilon_{\rm s}$ values.
@@
Finally, we put efforts to estimate the maximum mass outflow rate ($R^{\rm max}_{\dot{\rm m}}$) from the magnetized accretion disk. To achieve this, we compute $R_{\dot{\rm m}}^{\rm max}$ as a function of $\Upsilon_{\rm s}$ for two extreme cases of BH spin: a non-rotating black hole ($a_{\rm k}=0.0$) and a rapidly rotating black hole ($a_{\rm k}=0.99$). Here, we set $\alpha_{\rm B}=0.01$ and fix the plasma-$\beta$ at the inner critical point ($x_{\rm in}$) as $\beta_{\rm in}=100$, while allowing the energy ($\mathcal{E}_{\rm in}$) and angular momentum ($\lambda_{\rm in}$) at $x_{\rm in}$ to vary freely. The obtained results are presented in Fig. 4, where circles connected by solid (red) and dashed (blue) lines representing results for non-rotating ($a_{\rm k}=0.0$) and rapidly rotating ($a_{\rm k}=0.99$) black holes, respectively. Our findings clearly reveal that higher $a_{\rm k}$ drives stronger outflows regardless of the conduction parameter ($\Upsilon_{\rm s}$), which remain consistent with the results reported earlier in weak thermal conduction limit Jana and Das (2024). Interestingly, for a fixed $a_{\rm k}$, $R_{\dot{\rm m}}^{\rm max}$ initially increases with $\Upsilon_{\rm s}$, however eventually tends to saturate. Furthermore, we note that for a rapidly rotating BH with $a_{\rm k}=0.99$, $R_{\dot{\rm m}}^{\rm max}$ peaks at an impressive $25\%$. These results indeed suggest that accretion flows around rapidly spinning black holes with higher thermal conduction are significantly more susceptible to mass loss.